# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ 5.3 The OU regime near a nondegenerate critical point
Theorem 5.6 (OU linearization).
Under Assumption 5.5 and the conditions of Theorem 5.4, the diffusion (27) admits the local expansion
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s}+\rho_{1}(\Delta_{s})\,ds+\sqrt{\tau}\,\rho_{2}(\Delta_{s})\,dW_{s},$ (28)
where $\|\rho_{1}(\Delta)\|=o(\|\Delta\|)$ and $\|\rho_{2}(\Delta)\|_{\operatorname{op}}=o(1)$ as $\Delta\to 0$. Neglecting the higher-order remainder terms yields the OU SDE
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (29)
@@
A Taylor expansion gives $h(\theta^{\star}+\Delta)=H^{\star}\Delta+r(\Delta)$ with $\|r(\Delta)\|=o(\|\Delta\|)$. Continuity of $G^{\star}$ implies we can choose a continuous square root $C^{\star}(\theta)$ locally (e.g., the symmetric PSD square root), so
$C^{\star}(\theta^{\star}+\Delta)=C^{\star}(\theta^{\star})+R_{C}(\Delta),\qquad\|R_{C}(\Delta)\|_{\operatorname{op}}=o(1)\quad(\Delta\to 0).$
Substituting into (27) and collecting terms yields (28) with $\rho_{1}(\Delta):=-r(\Delta)$ and $\rho_{2}(\Delta):=R_{C}(\Delta)$. ∎
Corollary 5.7 (Stationary covariance and the Lyapunov equation).
If $H^{\star}\succ 0$, the OU process (29) is ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Since $H^{\star}\succ 0$, the matrix $H^{\star}$ is Hurwitz, so the continuous Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}=\tau\,G^{\star}(\theta^{\star})$ (30)
admits a unique symmetric positive definite solution $\Sigma_{\infty}$. In particular, the stationary covariance scales linearly with temperature $\tau=\eta/b$.
Because $H^{\star}$ is symmetric positive definite, left-multiplying (30) by $(H^{\star})^{-1}$ and taking traces gives the trace identity
$\mathrm{Tr}(\Sigma_{\infty})+\mathrm{Tr}\!\bigl((H^{\star})^{-1}\Sigma_{\infty}H^{\star}\bigr)=\tau\,\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr),$ (31)
which reduces to $2\,\mathrm{Tr}(\Sigma_{\infty})=\tau\,\mathrm{Tr}(G^{\star}H^{\star-1})$ when $H^{\star}$ and $\Sigma_{\infty}$ commute (e.g., when the eigenbases of curvature and noise align).
@@
Write $B:=\sqrt{\tau}\,(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=\tau G^{\star}(\theta^{\star})$. Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Apply Itô’s formula to $\Delta_{s}\Delta_{s}^{\top}$: the drift contributes $(H^{\star}\Sigma_{s}+\Sigma_{s}H^{\star\top})$ and the quadratic variation contributes $BB^{\top}$, giving $\frac{d}{ds}\Sigma_{s}=-H^{\star}\Sigma_{s}-\Sigma_{s}H^{\star\top}+BB^{\top}$. Setting $\frac{d}{ds}\Sigma_{s}=0$ yields (30). Uniqueness follows because for Hurwitz $H^{\star}$ the integral $\Sigma_{\infty}=\int_{0}^{\infty}e^{-H^{\star}s}\,(\tau G^{\star})\,e^{-H^{\star\top}s}\,ds$ converges and is the unique solution (Lemma A.3). ∎
Remark 5.8 (Likelihood case (Fisher-aligned diffusion)).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ and the OU diffusion reads
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C_{F}^{\star}\,dW_{s},\qquad C_{F}^{\star}{C_{F}^{\star}}^{\top}=F^{\star}(\theta^{\star}),$
making Fisher geometry explicit in both transient fluctuations and in the stationary Lyapunov balance (30).
@@ 5.4 Discrete-time stationary covariance: an exact bridge
The OU viewpoint is not the only consequence of the covariance identification. The same $G^{\star}$ governs the exact linearized discrete-time stationary covariance; the continuous-time Lyapunov equation emerges as the small-stepsize asymptotic reduction. This subsection makes the bridge explicit, so that the paper’s conclusions do not rest on the diffusion approximation alone.
@@
where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1})$ as in (25). If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), then the iterates are geometrically ergodic and the stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfies the discrete Lyapunov equation
@@
up to the same local remainder terms already used in the diffusion analysis.
@@
From (32), compute $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$, using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$. Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the iteration converges to the unique fixed point (33). The spectral radius condition $\rho(I-\eta H^{\star})<1$ is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$. ∎
Remark 5.10 (Continuous-time reduction and discrete-time corrections).
Expanding (33) and dividing by $\eta$, one recovers $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=\frac{\eta}{b}G^{\star}+\eta H^{\star}\Sigma_{\eta}H^{\star\top}$. Sending $\eta\downarrow 0$ with $\tau=\eta/b$ fixed drops the $O(\eta)$ correction term and yields the continuous Lyapunov equation (30). The correction is governed by $\eta\lambda_{\max}(H^{\star})$: when this product is small the two benchmarks nearly coincide; for moderate stepsizes the discrete equation is the correct comparator (Section 8). The covariance identification $\mathrm{Cov}(\xi_{t})\approx b^{-1}G^{\star}$ is logically prior to both analyses: the discrete Lyapunov equation is its exact linearized consequence, and the OU equation is the small-$\eta$ reduction.
@@ 6 Convergence Rates in Fisher Geometry
This section establishes matching upper and lower bounds for SGD when the error is measured in a frozen Fisher metric at a local optimum $\theta^{\star}$. Once the mini-batch noise covariance is identified (Section 4), the remaining work is a localized stochastic-approximation argument in a fixed geometry, followed by an information-theoretic lower bound expressed in the same metric. Convention. All rates in this section are stated in Fisher/Godambe risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
@@ 6.2 Upper Bound
Take total expectations in (37) to obtain $\mathbb{E}[V_{t+1}]\leq a_{t}\,\mathbb{E}[V_{t}]+u_{t}$ with $a_{t}:=1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}$ and $u_{t}:=\eta_{t}^{2}\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\eta_{t}^{2}\frac{c_{1}}{b}\mathrm{Tr}(F^{\star})\rho_{t}$. For $\eta_{t}=\eta_{0}/t$, there exists $t_{0}$ such that for $t\geq t_{0}$, $a_{t}\leq 1-\mu_{F}\eta_{0}/t$. Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$. Unrolling gives $\mathbb{E}[V_{T}]\leq A_{1,T}V_{1}+\sum_{t=1}^{T-1}A_{t+1,T}u_{t}$. Standard product bounds yield $A_{1,T}\leq CT^{-2\mu_{F}\eta_{0}}$ and $A_{t+1,T}\leq C(t/T)^{\mu_{F}\eta_{0}}$. Then, since $\mu_{F}\eta_{0}>1/2$, $\sum_{t\leq T}A_{t+1,T}\eta_{t}^{2}=O(1/T)$, giving the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term. The remainder term is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, yielding an additional $O((\mathrm{Tr}(F^{\star})/b)(1/T))$ contribution. This proves (38). The reduction to (6.5) follows from $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\mathrm{Tr}(F^{\star})$. ∎
@@ 6.3 Lower Bound
The lower bound proceeds in three layers of increasing generality. We first state the cleanest result in the parametric likelihood/Fisher case, then interpret it through oracle information, and finally discuss the extension to general losses.
6.3.1 Layer 1: Parametric Fisher-case lower bound.
@@
Theorem 6.8 (Fisher-case lower bound via van Trees).
Suppose Assumption 6.7 holds and the estimator $\hat{\theta}_{N}$ is based on $N$ i.i.d. observations from $P_{\theta^{\star}}$. Then
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (40)
for a universal constant $c>0$.
@@
This is the special case of Proposition 6.10 below with i.i.d. observations (so the oracle information bound holds with equality). We give the self-contained argument in the proof of Proposition 6.10. ∎
6.3.2 Layer 2: Oracle-information interpretation.
The heart of the lower bound is not i.i.d. sampling per se, but bounded predictable information growth. We now restate the result in terms of an abstract oracle transcript.
Assumption 6.9 (Oracle–information structure).
The oracle produces observations $Y_{t}=\psi(\theta_{t};Z_{t})$, where:
  1. (i)
    $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, with filtration $\mathcal{F}_{t}:=\sigma(Y_{1},\dots,Y_{t})$;
  2. (ii)
    Conditional unbiasedness: $\mathbb{E}_{\theta^{\star}}[Y_{t}\mid\mathcal{F}_{t-1}]=0$;
  3. (iii)
    Predictable information bound: $\mathbb{E}_{\theta^{\star}}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq\mathcal{I}_{1}(\theta^{\star})$, a.s., where $\mathcal{I}_{1}(\theta^{\star}):=F(\theta^{\star})=F^{\star}$ in the likelihood case;
  4. (iv)
    Uniform moment control: $\sup_{t}\mathbb{E}_{\theta^{\star}}\|Y_{t}\|^{4}<\infty$.
Under Assumption 6.9, the predictable quadratic variation of the oracle transcript satisfies $\sum_{t=1}^{N}\mathbb{E}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq N\mathcal{I}_{1}(\theta^{\star})$, so the total information available to any $\mathcal{F}_{N}$-measurable estimator is bounded by $N\mathcal{I}_{1}$.
Proposition 6.10 (Lower bound under predictable information budget).
Suppose Assumption 6.7 holds for the parametric family $\{P_{\theta}\}$ and the oracle satisfies Assumption 6.9 with $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$. Additionally, assume that each conditional law of $Z_{t}$ given $\mathcal{F}_{t-1}$ has a density differentiable in $\theta$, so that the conditional Fisher information $I_{t}(\theta\mid\mathcal{F}_{t-1})$ exists and satisfies $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$ a.s. Then every $\mathcal{F}_{N}$-measurable estimator $\hat{\theta}_{N}$ satisfies
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (41)
for a universal constant $c>0$.
In particular, if one SGD iteration uses a mini-batch of size $b$ (so $N=Tb$ oracle calls), then $\inf_{\hat{\theta}_{T}}\sup_{\theta}\mathbb{E}[\|\hat{\theta}_{T}-\theta\|_{F(\theta)}^{2}]\geq c\,d/(Tb)$.
Proof.
The argument extends Theorem 6.8 from i.i.d. observations to the adaptive oracle transcript. The van Trees bound requires an upper bound on the Fisher information $I_{N}(\theta)$ of the joint likelihood of $(Y_{1},\ldots,Y_{N})$ with respect to $\theta$. Three ingredients deliver this bound.
Step 1 (Adaptivity does not inflate per-step information). At step $t$, the query $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, but $Z_{t}\sim P_{\theta}$ is drawn independently of $\mathcal{F}_{t-1}$ given $\theta$. Hence the conditional distribution of $Z_{t}$ given $\mathcal{F}_{t-1}$ depends on $\theta$ only through $\{P_{\theta}\}$, not through the past transcript.
Step 2 (Second-moment bound == per-step Fisher bound). Under correct specification, $Y_{t}=-s_{\theta}(Z_{t})$ is the negative score. The conditional Fisher information of $Z_{t}$ given $\mathcal{F}_{t-1}$ about $\theta$ is $I_{t}(\theta\mid\mathcal{F}_{t-1})=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}\mid\mathcal{F}_{t-1}]=F(\theta)$—exactly the score second-moment matrix. So Assumption 6.9(iii) directly bounds $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$. To extend from $\theta^{\star}$ to all $\theta\in\Theta_{\operatorname{loc}}$, continuity of $F(\theta)$ (Assumption 6.7(ii)) gives $F(\theta)\preceq(1+\epsilon)F^{\star}$ on a sufficiently small $\Theta_{\operatorname{loc}}$; the factor $(1+\epsilon)$ is absorbed into the universal constant $c$.
Step 3 (Chain rule and conclusion). The chain rule for conditional Fisher information applied to the sequential factorization of the joint likelihood gives
$I_{N}(\theta)=\sum_{t=1}^{N}\mathbb{E}_{\theta}\!\bigl[I_{t}(\theta\mid\mathcal{F}_{t-1})\bigr]\preceq N\,(1{+}\epsilon)\,F^{\star}.$
Substituting into van Trees with prior $\pi$ and sending $\epsilon\to 0$ yields (41). ∎
Remark 6.11 (Beyond i.i.d. oracles).
The classical i.i.d. setting satisfies Assumption 6.9 with equality in (iii). The assumption also accommodates exchangeable or conditionally i.i.d. sampling schemes (e.g., de Finetti mixtures) provided the per-step conditional information is bounded by $\mathcal{I}_{1}$. What is required is a predictable information bound on the oracle transcript, not independence across calls. The critical logical step is that the van Trees inequality applies to the joint transcript likelihood, whose total Fisher information is controlled by summing the per-step conditional bounds.
Remark 6.12 (What enters the lower bound).
The lower bound (41) involves $d$ (ambient dimension) and $N$ (oracle budget). The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ enter the upper bound and oracle complexity (Theorems 6.5 and 7.3) but not the lower bound, which is stated in terms of $d$ to give the sharpest minimax comparison. When $d_{\mathrm{eff}}\ll d$, the gap between upper and lower bounds tightens correspondingly.
6.3.3 Layer 3: General-loss / Godambe extension.
Remark 6.13 (Extension to general losses).
In the general-loss case, the same lower-bound form $\Omega(d/N)$ is recovered under local asymptotic conditions that identify the relevant sandwich geometry. Specifically, if the loss satisfies a local asymptotic normality (LAN) condition at $\theta^{\star}$ with information matrix $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$, then the van Trees argument carries through with $F^{\star}$ replaced by $G^{\star}(\theta^{\star})$ and the dual norm adjusted accordingly. The LAN condition is standard for smooth M-estimators under regularity conditions comparable to those already imposed in Section 3; see Le Cam (1986) and van der Vaart (1998). We state this extension as a remark rather than a theorem because the full verification of LAN for the adaptive oracle transcript of Assumption 6.9 requires additional technical conditions (uniform local asymptotic equicontinuity) that we do not develop here.
Corollary 6.14 (Matching rates (up to constants)).
Under Theorem 6.5 and the oracle model of Assumption 6.9,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]=\Theta\!\left(\frac{1}{Tb}\right)\quad\text{equivalently}\quad\Theta\!\left(\frac{1}{N}\right)\text{ per oracle call},\quad N=Tb.$

### 15 Apr 2026 v4

@@ 5.4 The OU regime as the equilibrium specialization
Theorem 5.10 (OU fluctuation limit at a stationary point).
Under Assumption 5.9 and the conditions of Theorem 5.7 with $\theta_{0}=\theta^{\star}$, the fluctuation limit $U$ of Theorem 5.7 solves the Ornstein–Uhlenbeck SDE
$dU_{t}\;=\;-H^{\star}\,U_{t}\,dt\;+\;C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (32)
Equivalently, on the original iterate scale, the centered iterate $\hat{\theta}^{\eta}(t)-\theta^{\star}$ admits the asymptotic expansion
$\hat{\theta}^{\eta}(t)-\theta^{\star}\;=\;\sqrt{\tfrac{\eta}{b}}\,U_{t}\;+\;o_{\mathbb{P}}\!\bigl(\sqrt{\tfrac{\eta}{b}}\bigr).$
@@
Specialize (26) to $\bar{\theta}(\cdot)\equiv\theta^{\star}$: the drift becomes $H^{\star}U_{t}$ and the diffusion becomes $C^{\star}(\theta^{\star})dW_{t}$. The original-scale expansion is (27) with $\bar{\theta}\equiv\theta^{\star}$. ∎
Corollary 5.11 (Fluctuation-scale Lyapunov equation and discrete-scale benchmark).
Under Assumption 5.9, the OU process (32) is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{U})$, where $\Sigma_{U}$ is the unique symmetric positive definite solution of the fluctuation-scale Lyapunov equation
$H^{\star}\,\Sigma_{U}\;+\;\Sigma_{U}\,(H^{\star})^{\top}\;=\;G^{\star}(\theta^{\star}).$ (33)
On the original iterate scale, Proposition 5.13 provides a discrete Lyapunov covariance $\Sigma_{\eta}^{\mathrm{lin}}$ for the linearized recursion, which at leading order in $\eta$ equals $(\eta/b)\Sigma_{U}$ and satisfies
$H^{\star}\,\Sigma_{\eta}^{\mathrm{lin}}+\Sigma_{\eta}^{\mathrm{lin}}\,(H^{\star})^{\top}\;=\;\frac{\eta}{b}\,G^{\star}(\theta^{\star})+O(\eta)\cdot\Sigma_{\eta}^{\mathrm{lin}}.$ (34)
This is the linearized-recursion benchmark on the original iterate scale; we do not claim an invariant law for the nonlinear SGD iterates here. Because $H^{\star}$ is symmetric positive definite, cyclicity of trace gives the unconditional trace identity
$2\,\mathrm{Tr}(\Sigma_{U})\;=\;\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr).$ (35)
@@
Write $B:=(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=G^{\star}(\theta^{\star})$. Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{U})$. Applying Itô’s formula to $U_{t}U_{t}^{\top}$ and taking stationary expectations yields $(H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top})+BB^{\top}=0$, which is (33); the integral representation $\Sigma_{U}=\int_{0}^{\infty}e^{-H^{\star}s}G^{\star}(\theta^{\star})e^{-H^{\star\top}s}ds$ converges and gives uniqueness (Lemma A.1). Equation (34) follows from Proposition 5.13 after dividing out leading powers of $\eta$ (Remark 5.14). The trace identity follows by left-multiplying (33) by $(H^{\star})^{-1}$ and using $\mathrm{Tr}(H^{-1}\Sigma H)=\mathrm{Tr}(\Sigma)$. ∎
Remark 5.12 (Likelihood specialization).
When $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$. The fluctuation-scale Lyapunov equation (33) then reads $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=F^{\star}(\theta^{\star})$; the linearized discrete-recursion benchmark of Proposition 5.13 satisfies $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ at leading order in $\eta$ (Corollary 5.11). Fisher geometry is therefore explicit both in the transient fluctuations and in the linearized stationary Lyapunov balance; we do not assert an invariant law for the nonlinear SGD iterates here.
@@ 5.5 Discrete-time stationary covariance: an exact bridge
The same $G^{\star}$ that governs the fluctuation-scale Lyapunov equation (33) also governs the exact linearized discrete-time stationary covariance. The continuous-time identity (34) is the small-$\eta$ reduction of the discrete Lyapunov equation below.
@@
where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})$ as in (24). If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), the linearized covariance recursion is stable and admits a unique stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfying the discrete Lyapunov equation
@@
up to the same local remainder terms already used in the fluctuation analysis.
@@
From (36), $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$ using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$. Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the covariance iteration converges to the unique fixed point (37). The spectral radius condition is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$. The statement is about the second-moment recursion for the linearized iterates; geometric ergodicity of the nonlinear SGD iterates would require additional assumptions (drift + minorization) that we do not impose. ∎
Remark 5.14 (Small-stepsize reduction and the continuous limit).
Expanding (37), $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=(\eta/b)G^{\star}+\eta\,H^{\star}\Sigma_{\eta}H^{\star\top}$. When $\eta\lambda_{\max}(H^{\star})$ is small, the $O(\eta)$ discretization correction is negligible, and the discrete equation reduces to the continuous-time, original-scale identity (34). The covariance identification $\mathbb{E}[\xi\xi^{\top}\mid\mathcal{F}]\approx b^{-1}G^{\star}$ is logically prior to both: the discrete Lyapunov equation is its exact linearized consequence, and the continuous equation is the small-$\eta$ reduction that matches the fluctuation-scale limit of Theorem 5.7 after multiplying back by $\eta/b$.
@@ 6 Convergence Rates in Fisher Geometry
This section establishes matching upper and lower bounds for SGD when the error is measured in a frozen Fisher metric at a local optimum $\theta^{\star}$. Once the mini-batch noise covariance is identified (Section 4), the remaining work is a localized stochastic-approximation argument in a fixed geometry, followed by an information-theoretic lower bound expressed in the same metric. Convention. All rates in this section are stated in the identified statistical risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
@@ 6.2 Upper Bound
Take total expectations in (41) to obtain $\mathbb{E}[V_{t+1}]\leq a_{t}\,\mathbb{E}[V_{t}]+u_{t}$ with $a_{t}:=1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}$ and $u_{t}:=\eta_{t}^{2}\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\eta_{t}^{2}\frac{c_{1}}{b}\mathrm{Tr}(F^{\star})\rho_{t}$. Fix any $\epsilon\in(0,2\mu_{F})$. For $\eta_{t}=\eta_{0}/t$ and $t$ large enough that $c_{0}\eta_{0}^{2}/t\leq\epsilon\eta_{0}$, we have $a_{t}\leq 1-(2\mu_{F}-\epsilon)\eta_{0}/t$; set $\alpha:=(2\mu_{F}-\epsilon)\eta_{0}$, which satisfies $\alpha>1$ whenever $\eta_{0}>1/(2\mu_{F})$ and $\epsilon$ is chosen small enough. Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$ (Lemma A.2). Unrolling gives $\mathbb{E}[V_{T}]\leq A_{1,T}V_{1}+\sum_{t=1}^{T-1}A_{t+1,T}u_{t}$. With the consistent exponent $\alpha$, the product bounds are $A_{1,T}\leq C\,T^{-\alpha}$ and $A_{t+1,T}\leq C\,(t/T)^{\alpha}$. Since $\alpha>1$, the standard Robbins–Monro calculation gives $\sum_{t=1}^{T-1}A_{t+1,T}\,\eta_{t}^{2}=O(1/T)$, yielding the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term. The remainder contribution is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, adding $O((\mathrm{Tr}(F^{\star})/b)(1/T))$. The transient term $A_{1,T}V_{1}=O(T^{-\alpha})$ appears in the bound (42) with exponent $\alpha\leq 2\mu_{F}\eta_{0}$; sending $\epsilon\downarrow 0$ gives the exponent $2\mu_{F}\eta_{0}$ stated in (42) up to an arbitrarily small loss. The reduction to (6.5) follows from $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\mathrm{Tr}(F^{\star})$. ∎
@@ 6.3 Lower Bound
We state the clean lower bound in the local parametric Fisher setting with i.i.d. observations. Extensions to adaptive oracles and general losses require additional machinery and are left as discussion (Remarks 6.116.12).
@@
Proposition 6.8 (Local i.i.d. Fisher-case lower bound via van Trees).
Suppose Assumption 6.7 holds. Consider the local estimation model in which, for each $\theta\in\Theta_{\operatorname{loc}}$, one observes $N$ i.i.d. samples from $P_{\theta}$ and forms an estimator $\hat{\theta}_{N}$ (a measurable function of the samples). Then
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c_{\mathrm{loc}}\,d}{N},$ (44)
for a constant $c_{\mathrm{loc}}>0$ that depends on the local regularity of $F(\theta)$ on $\Theta_{\operatorname{loc}}$ and on the prior regularity $J_{\pi}$ from Assumption 6.7(iv). In particular, if each SGD iteration uses a fresh mini-batch of size $b$ (so $N=Tb$ fresh-sample gradient calls), the Fisher-metric risk is bounded below by $c_{\mathrm{loc}}\,d/(Tb)$ on the i.i.d. population oracle.
@@
Standard van Trees argument (Gill and Levit, 1995). By Assumption 6.7(ii), the per-coordinate local Fisher scale $\bar{f}:=\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathrm{Tr}F(\theta)/d<\infty$ is $N$-independent. Fix a smooth prior $\pi$ on $\Theta_{\operatorname{loc}}$ with finite $J_{\pi}$ (Assumption 6.7(iv)); the van Trees inequality for the Fisher-weighted Bayes risk gives $r_{\pi}(\hat{\theta}_{N})\geq d/(N\bar{f}+J_{\pi}/d)$, and for $N\geq N_{0}:=J_{\pi}/(d\bar{f})$ the denominator is at most $2N\bar{f}$, yielding $r_{\pi}(\hat{\theta}_{N})\geq c_{\mathrm{loc}}\,d/N$ with $c_{\mathrm{loc}}:=1/(2\bar{f})>0$ depending only on the local regularity of $F(\theta)$ on $\Theta_{\operatorname{loc}}$ and independent of both $N$ and $d$. Minimax risk dominates Bayes risk for any such prior, proving (44). ∎
Remark 6.9 (Noise-covariance condition used in the upper bounds).
The upper bounds of Theorem 6.5 and the oracle complexity of Theorem 7.3 use the conditional-covariance bound $\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\preceq b^{-1}F^{\star}$, which is exactly the fresh-sampling specialization of Theorem 4.3 in the correctly specified likelihood case. No structure beyond this covariance identification is required.
Remark 6.10 (Rate matching, not geometry matching).
Proposition 6.8 establishes a lower bound of rate order $\Omega(d/N)$ in the i.i.d. parametric Fisher setting. The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ appear only in the upper bound and oracle complexity (Theorems 6.5 and 7.3), not in the lower bound. Accordingly, we claim rate-order matching in the local Fisher case; we do not claim that the lower bound matches the geometry dependence of the upper bound.
Remark 6.11 (Adaptive-oracle and general-loss extensions).
Two natural extensions of Proposition 6.8 are left to future work. (a) Adaptive oracles. A sequential oracle transcript $Y_{t}=\psi(\theta_{t};Z_{t})$ with a predictable per-step Fisher-information bound should yield (44) through a sequential van Trees argument, provided one establishes a uniform chain-rule bound on the joint transcript likelihood (requiring regularity conditions such as uniform local asymptotic equicontinuity that we do not develop here). (b) General losses. A local asymptotic normality (LAN) condition at $\theta^{\star}$ with Godambe information $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$ would produce a sandwich analogue of (44); see Le Cam (1986) and van der Vaart (1998).
Remark 6.12 (Legacy labels for cross-references).
These labels are retained for backward compatibility; the i.i.d. Fisher lower bound is Proposition 6.8, and the general-loss extension is discussed in Remark 6.11.
Corollary 6.13 (Matching rate order in the local Fisher case).
Under Theorem 6.5 and the noise identification of Theorem 4.3 in the correctly specified likelihood setting,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]\;=\;\Theta\!\left(\frac{1}{N}\right),\qquad N=Tb,$
at the level of rate order. Dependence on intrinsic geometry ($d_{\mathrm{eff}}$, $\kappa_{F}$) appears only in the upper bound.